# Probing the Bounce Energy Scale in Bouncing Cosmologies with Pulsar Timing Arrays

### v1

@@ Abstract
In this work we constrain the bounce energy scale $\rho_{s\downarrow}^{1/4}$ in a generic framework of bouncing cosmologies using the nanohertz stochastic gravitational-wave background recently detected by pulsar-timing arrays (NANOGrav 15-yr, EPTA DR2, PPTA DR3, IPTA DR2). A full Bayesian fit of the analytic SGWB spectrum for this bounce scenario reveals, for the first time, two distinct posterior branches in $(\rho_{s\downarrow}^{1/4},w_{1})$: one near $w_{1}\approx 0.3$ and one at $w_{1}\gg 1$, where $w_{1}$ is the contraction-phase equation-of-state. We find that the bouncing model attains larger Bayes factors against each of six conventional SGWB sources (SMBHBs, inflationary GWs, cosmic strings, domain walls, first-order phase transitions, scalar-induced GWs), demonstrating strong preference of current PTA data for the bounce hypothesis. Compared to the more generic “dual inflation–bounce” scenario, the concrete bounce realization yields smaller Bayes factors, indicating that PTA measurements impose tighter constraints when the bounce scale is explicit. Moreover, the two posterior branches illuminate distinct theoretical frontiers. The “right” branch ($w_{1}\gg 1$) violates the dominant energy condition (DEC), thereby providing direct empirical impetus for models with novel early-Universe physics—e.g. ghost condensates, higher-derivative or modified-gravity operators, and extra-dimensional effects. Independently, both branches infer $\rho_{s\downarrow}^{1/4}$ above the Planck scale $M_{\rm pl}$, demonstrating that current PTAs already probe trans-Planckian regimes. Together, these findings offer a rare observational window into UV completions of cosmology. We further describe how normalizing-flow–based machine-learning can accelerate such Bayesian analyses as PTA data volumes increase.
I Introduction
@@
Currently, multiple pulsar timing array (PTA) collaborations—including NANOGrav [1, 2], EPTA [3, 4], PPTA [5, 6], IPTA [7], and CPTA [8]—report a common nanohertz‐frequency stochastic gravitational wave background (SGWB) with Hellings–Downs correlations detected at up to $3-4\sigma$ significance. This breakthrough opens a unique observational window onto both astrophysical and cosmological gravitational wave (GW) sources [9, 10, 11, 12, 13, 14], including supermassive–black‐hole binaries [15, 16, 17, 18], inflationary GWs [19, 20, 21], cosmic strings [22, 23, 24], domain walls [25, 26, 27], first‐order phase transitions [28, 29, 30, 31, 32, 33], scalar‐induced GWs [34, 35, 36], and more exotic early‐Universe scenarios such as audible axions [37, 38], kination domination [39, 40], preheating [41, 42], and particle‐production–driven signals [43, 44]. Together, these results inaugurate a powerful new probe of fundamental physics in cosmology and gravity.
@@
Remarkably, this nanohertz SGWB signal also provides a novel test of bouncing cosmologies [45] (see review in [14] and recent developments in [46, 47, 48, 49]). Unlike standard inflationary scenarios—which posit a singular origin [50, 51, 52, 53, 54, 55, 56, 57] (cf. recent discussions in [58, 59])—bouncing models avoid the singularity by passing through a finite, nonzero minimum scale factor during a contraction–expansion transition [60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89] (see reviews in [90, 91, 92, 93]). Bayesian comparisons between generic inflation and bounce scenarios yield Bayes factors near unity—evidence for an inflation–bounce duality in SGWB data [45, 94] (see Eq. (8) and Table 3).
@@
Previous Bayesian treatments of bouncing models in PTA analyses [45] did not include the explicit bounce energy scale $\rho_{s\downarrow}^{1/4}$, since the full evolution of primordial gravitational waves through the non-singular bounce was not computed. As illustrated in the right panel of Fig. 1, the scale factor of universe attains its minimum—and thus the quasi-highest energy scale—at the bounce point. Incorporating $\rho_{s\downarrow}^{1/4}$ therefore requires tracking the horizon crossings of PGWs across all four phases of the bounce, leading directly to the spectrum in Eq. (1). In particular, Figure 2 illustrates that $\rho_{s\downarrow}^{1/4}$ corresponds to the maximum Hubble scale reached immediately after the non-singular bounce (onset of Phase IV). Since additional microphysical effects could drive the instantaneous energy even higher at the bounce point, we refer to $\rho_{s\downarrow}^{1/4}$ as the quasi-highest energy scale of the bouncing universe.
@@
Figure 1: Evolution of the comoving Hubble radius $|aH|^{-1}$ in (left) inflationary and (right) bouncing cosmologies [95, 96, 97]. In the bounce the horizon shrinks in Phases I and III and grows in Phases II and IV; primordial GWs exit in I, reenter in II, exit again in III, and finally reenter in IV, producing a double horizon–crossing history [96]. Inflation maintains an almost constant “inflationary” energy scale, tightly constrained by current GW limits [14]. By contrast, bouncing models attain their maximum (“bounce”) energy scale at the central bounce [96, 97], which we constrain here using PTA data.
@@
In a recent work [96] we established a general framework for the generation and evolution of primordial gravitational waves (PGWs) in non-singular bouncing cosmologies. We partition the cosmic history into four phases—contracting collapse (I), bouncing contraction (II), bouncing expansion (III), and post-bounce radiation (IV)—as illustrated in Fig. 1. Employing the GW–propagation–matrix method of Ref. [96], we derived an analytical expression for the PGW spectrum in bouncing cosmologies (c.f. Eq. (53) of Ref. [96]).
Building on this, Ref. [97] specialized to a concrete bounce by imposing three assumptions: (i) Phase IV is radiation dominated ($w_{4}=1/3$), as in the standard post-reheating Universe; (ii) the immediate pre- and post-bounce phases are governed by an exotic fluid with $w_{2}=w_{3}\to-\infty$ (e.g. Quintom matter [98]), ensuring a non-singular bounce; and (iii) the bounce is symmetric in conformal time, $\eta_{s\downarrow}\equiv\eta_{1\downarrow}=\eta_{3\downarrow}$. Under these conditions one obtains the closed-form SGWB spectrum in Eq. (1) (c.f. Eq. (24) of Ref. [97]), which explicitly depends on the bounce energy scale $\rho_{s\downarrow}^{1/4}$ (Eq. (2)).
In this work we employ the explicit SGWB spectrum for a non-singular bounce (Eq. (1)) to place the first direct PTA constraints on the bounce energy scale $\rho_{s\downarrow}^{1/4}$ in a generic four-phase bouncing-cosmology framework. Fitting to the combined NANOGrav 15-yr, EPTA DR2, PPTA DR3 and IPTA DR2 datasets [2, 4, 6, 7, 99, 100, 101, 102, 103, 104, 105, 106] via full Bayesian inference, we uncover two distinct posterior branches in $(\rho_{s\downarrow}^{1/4},w_{1})$: one at $w_{1}\approx 0.3$ and one at $w_{1}\gg 1$. Recasting these bounds into the dual $(a_{s\downarrow},w_{1})$ parametrization yields fully consistent results.
@@
    Dominant-energy‐condition violation. The “right” branch ($w_{1}\gg 1$) violates the dominant energy condition (Table 1), offering empirical impetus to explore novel pre-bounce physics such as ghost condensates, higher-derivative or modified-gravity operators, and extra-dimensional effects.
@@
    consistent with the lower bound found in Ref. [97]. Thus, current PTAs are already probing trans-Planckian regimes in bouncing models. Any proposed UV completion must respect these observational constraints, providing a rare window into physics beyond the Planck scale.
@@
II Stochastic Gravitational Waves for Bouncing Cosmologies
In Ref. [97], considering a realistic bouncing cosmology that reheats to the standard radiation dominated era ($w_{4}=1/3$) after a sharp, symmetric bounce ($w_{2}=w_{3}\to-\infty$, $\eta_{s\downarrow}\equiv\eta_{1\downarrow}=\eta_{3\downarrow}$), we derive the PGW‑induced stochastic gravitational‑wave background (SGWB) spectrum (c.f. Eq.(24) in Ref. [97]), which encodes the bounce energy scale $\rho_{s\downarrow}^{1/4}$:
$\Omega_{\mathrm{GW}}(f)\,h^{2}=\frac{h^{2}}{24}\left(\frac{f_{H_{0}}}{f_{m_{\mathrm{pl}}}}\right)^{2}\frac{C(w_{1})}{(2\pi)^{-n_{T}(w_{1})-1}}\left(\frac{f}{f_{H_{0}}}\right)^{n_{T}(w_{1})}\left[\frac{\rho_{s\downarrow}^{1/4}}{(\rho_{c0}/\Omega_{\gamma 0})^{1/4}}\right]^{4-n_{T}(w_{1})}\mathcal{T}_{\mathrm{eq}}(f)\,.$ (1)
Here, $\rho_{s\downarrow}^{1/4}$ is defined at the beginning of Phase IV (the reheating point) with the assumption that Phase IV is standard post-reheating radiation-dominated era (c.f. Eq.(20) in Ref. [97]),
$\rho_{s\downarrow}\equiv\rho(\eta_{s})=3H^{2}(\eta_{s\downarrow})m_{\mathrm{pl}}^{2}~{},$ (2)
$C(w_{1})$ is the amplitude factor [97],
@@ II Stochastic Gravitational Waves for Bouncing Cosmologies
$n_{T}(w_{1})$ is the tilt of the primordial tensor spectrum, depending on $w_{1}$ [97],
$n_{T}\equiv\frac{\mathrm{d}\ln\mathcal{P}_{h}(\eta_{k})}{\mathrm{d}\ln k}=\begin{cases}5-2x(w_{1}),&-\tfrac{1}{3}\leq w_{1}<1.\\ 3+2x(w_{1}),&w_{1}\geq 1,\end{cases},$ (4)
where $w_{i}$ denotes the equation‑of‑state parameter in phase $i$, $x(w_{1})\equiv\tfrac{3(1-w_{1})}{2(3w_{1}+1)}$ and $\Gamma(z)=\int_{0}^{\infty}t^{z-1}e^{-t}\,\mathrm{d}t$ is the Gamma function. The present‑day critical density is $\rho_{c0}=3H_{0}^{2}m_{\mathrm{pl}}^{2}=8.098\,h^{2}\times 10^{-11}\,\mathrm{eV}^{4}$ with the reduced Hubble parameter $h=0.677$. Reference frequencies $f_{H_{0}}=2.2\times 10^{-18}\,\mathrm{Hz}$ and $f_{m_{\mathrm{pl}}}=3.7\times 10^{42}\,\mathrm{Hz}$ correspond to $H_{0}$ and the reduced Planck mass $m_{\mathrm{pl}}\equiv(\sqrt{8\pi G})^{-1}$, respectively, in natural units ($\hbar=c=1$). The radiation density fraction today is $\Omega_{\gamma 0}h^{2}=2.474\times 10^{-5}$. The transfer function
@@
captures the evolution from radiation to matter domination, with $f_{\mathrm{eq}}=2.01\times 10^{-17}\,\mathrm{Hz}$ marking matter–radiation equality. For the frequency band accessible to present gravitational‑wave experiments—CMB/BICEP, pulsar‑timing arrays, and Advanced LIGO–Virgo—the equality‑era transfer function satisfies $\mathcal{T}_{\mathrm{eq}}(f)\simeq 1$ because $f\gg f_{\mathrm{eq}}$. Accordingly, the stochastic gravitational‑wave background observed today inherits the primordial tensor tilt $n_{T}(w_{1})$.
Figure 3 shows the tensor spectral tilt $n_{T}(w_{1})$ (red curve) obtained from Eq. (4). The stochastic background is red‑tilted ($n_{T}<0$) for $w_{1}<0$, blue‑tilted ($n_{T}>0$) for $w_{1}>0$, and strictly scale‑invariant ($n_{T}=0$) at $w_{1}=0$ (a matter‑dominated bounce). Consequently, low‑frequency detectors (CMB/BICEP)—most sensitive to red tilts—set the strongest bounds in the $w_{1}<0$ regime, whereas high‑frequency instruments (aLIGO+aVirgo) constrain more tightly when $w_{1}>0$.
@@
Current pulsar‑timing‑array (PTA) measurements prefer a blue-tilted spectrum: the NANOGrav 15‑year data yield $n_{T}=1.8\pm 0.3$ (68% C.L.) [107], shown as the yellow band in Fig. 3. Its intersection with the red curve selects two allowed intervals,
@@
III Interpretation of Pulsar-Timing-Array Data in Bouncing Cosmology
III.1 Posterior distribution of bounce energy scale and physics interpretation
To determine the region in the joint parameter space $(\rho_{s\downarrow}^{1/4},w_{1})$ of bouncing cosmologies that is compatible with current PTA measurements, we carry out a Bayesian fit of Eq. (1) to the full PTA dataset ( NANOGrav 15-year, EPTA DR2, PPTA DR3 and IPTA DR2) [2, 4, 6, 7, 99, 100, 101, 102, 103, 104, 105, 106]. In Fig. 4 we present the resulting two-dimensional posterior distribution in the $(\rho_{s\downarrow}^{1/4},w_{1})$ plane for the bouncing scenario. Fig. 4 shows two distinct posterior branches, one near $w_{1}\approx 0.3$ and the other at $w_{1}\gg 1$, consistent with our analytic behavior of $n(w_{1})$ in Eq. (6). The small offset between the posterior peaks and the analytic prediction ($0.370.20<w_{1}<$ 0.37 and $w_{1}>13$) arises from the statistical uncertainties in the PTA data. More precisely, Fig. 4 shows that on the “left” branch the inferred bounce scale $\rho_{s\downarrow}^{1/4}$ grows with $w_{1}$ (see the high-resolution inset in Fig. 5), reflecting the monotonic rise of $n_{T}(w_{1})$ in that region, whereas on the “right” branch $\rho_{s\downarrow}^{1/4}$ asymptotes to a constant value as $n_{T}(w_{1})\to 2$ for $w_{1}\gg 1$ (the amplitude factor $C(w_{1})$ must also be accounted for in a more detailed description). Additionally, in Fig. 6, we show violin plots for two representative points—one from each of the $1\sigma$ posterior branches in the $(\rho_{s\downarrow}^{1/4},w_{1})$ plane (Fig. 4)—alongside examples in the $(a_{s\downarrow},w_{1})$ parametrization (Fig. 8) and for six conventional SGWB source models (Sec. A). These plots confirm that both bounce-model branches identified in this work can explain the nanohertz SGWB observed by current PTA datasets.

### v4

@@ Abstract
In this work, we constrain the bounce energy scale $\rho_{s\downarrow}^{1/4}$ in a generic bouncing-cosmology framework using the probable nanohertz stochastic gravitational-wave signal from NANOGrav 15-yr, EPTA DR2, PPTA DR3, and IPTA DR2. A full Bayesian fit to the analytic stochastic-GW spectrum, for the first time, reveals two distinct posterior branches in $\bigl{(}\rho_{s\downarrow}^{1/4},\,w_{1}\bigr{)}$: one near $w_{1}\approx 0.3$ and another at $w_{1}\gg 1$, where $w_{1}$ is the contraction-phase equation of state. We find that current PTA data strongly favor the bounce hypothesis over six conventional SGWB sources, while yielding smaller Bayes factors than its parent dual inflation–bounce scenario as the bounce scale is explicit. Both branches imply $\rho_{s\downarrow}^{1/4}>M_{\rm pl}$, indicating a trans-Planckian probe. In particular, the $w_{1}\gg 1$ branch violates the dominant energy condition, motivating novel early-Universe physics. These results, confirmed by an accelerated a machine learning approach, offer a rare observational window into UV completions of cosmology.
journal: Eur. Phys. J. C
1 Introduction
@@
Currently, multiple pulsar timing array (PTA) collaborations—including NANOGrav NANOGrav:2020spf ; NANOGrav:2023gor , EPTA EPTA:2021crs ; EPTA:2023fyk , PPTA Goncharov:2021oub ; Reardon:2023gzh , IPTA Antoniadis:2022pcn , and CPTA Xu:2023wog —report a common-spectrum process with Hellings–Downs correlations detected at up to $3-4\sigma$ significance, which may be the first hint of a nanohertz stochastic gravitational wave background (SGWB). This breakthrough opens a unique observational window onto both astrophysical and cosmological gravitational wave (GW) sources NANOGrav:2023hvm ; EPTA:2023xxk ; Figueroa:2023zhu ; Bian:2023dnv ; Ellis:2023oxs ; Caprini:2018mtu , including supermassive–black-hole binaries Sesana:2013wja ; Kelley:2017lek ; Chen:2018znx ; Burke-Spolaor:2018bvk ; Calza:2024qxn , inflationary GWs Zhao:2013bba ; Guzzetti:2016mkm ; Vagnozzi:2020gtf , cosmic strings Siemens:2006yp ; Cui:2018rwi ; Gouttenoire:2019kij , domain walls Hiramatsu:2013qaa ; Ferreira:2022zzo ; Bian:2022qbh , first-order phase transitions Hindmarsh:2013xza ; Hindmarsh:2015qta ; Caprini:2015zlo ; Hindmarsh:2017gnf ; Gouttenoire:2023bqy ; Salvio:2023ynn , scalar-induced GWs Ananda:2006af ; Baumann:2007zm ; Kohri:2018awv , and more exotic early-Universe scenarios such as audible axions Machado:2019xuc ; Machado:2018nqk , kination domination Co:2021lkc ; Oikonomou:2023qfz , preheating Bethke:2013aba ; Adshead:2019igv , and particle-production–driven signals Dimastrogiovanni:2016fuu ; DAmico:2021fhz . Together, these results inaugurate a powerful new probe of fundamental physics in cosmology and gravity.
@@
Remarkably, this nanohertz SGWB signal also provides a novel test of bouncing cosmologies Li:2024oru (see review in Caprini:2018mtu and recent developments in Zhu:2023lbf ; Papanikolaou:2024fzf ; Ben-Dayan:2024aec ; Qiu:2024sdd ). Unlike standard inflationary scenarios—which posit a singular origin Guth:1980zm ; Starobinsky:1980te ; Sato:1980yn ; Linde:1981mu ; Albrecht:1982wi ; Mukhanov:1990me ; Borde:1993xh ; Borde:2001nh (cf. recent discussions in Lesnefsky:2022fen ; Easson:2024uxe )—bouncing models avoid the singularity by passing through a finite, nonzero minimum scale factor during a contraction–expansion transition Khoury:2001wf ; Gasperini:2002bn ; Creminelli:2006xe ; Peter:2006hx ; Cai:2007qw ; Cai:2008qw ; Saidov:2010wx ; Li:2011nj ; Cai:2011tc ; Easson:2011zy ; Bhattacharya:2013ut ; Qiu:2015nha ; Cai:2016hea ; Barrow:2017yqt ; deHaro:2017yll ; Ijjas:2018qbo ; Boruah:2018pvq ; Nojiri:2019yzg ; Silva:2015qna ; Silva:2020bnn ; Silva:2023ieb ; Nayeri:2005ck ; Brandenberger:2006xi ; Fischler:1998st ; Cai:2009rd ; Li:2014era ; Cheung:2014nxi ; Li:2014cba ; Li:2015egy ; Li:2020nah (see reviews in Novello:2008ra ; Brandenberger:2016vhg ; Nojiri:2017ncd ; Odintsov:2023weg ). Bayesian comparisons between generic inflation and bounce scenarios yield Bayes factors near unity—evidence for an inflation–bounce duality in SGWB data Li:2024oru ; Li:2013bha (see Eq. (8) and Table 3).
@@
Previous Bayesian treatments of bouncing models in PTA analyses Li:2024oru did not include the explicit bounce energy scale $\rho_{s\downarrow}^{1/4}$, since the full evolution of primordial gravitational waves through the non-singular bounce was not computed. As illustrated in the right panel of Fig. 1, the scale factor of universe attains its minimum—and thus the quasi-highest energy scale—at the bounce point. Incorporating $\rho_{s\downarrow}^{1/4}$ therefore requires tracking the horizon crossings of PGWs across all four phases of the bounce, leading directly to the spectrum in Eq. (1). In particular, Figure 2 illustrates that $\rho_{s\downarrow}^{1/4}$ corresponds to the maximum Hubble scale reached immediately after the non-singular bounce (onset of Phase IV). Since additional microphysical effects could drive the instantaneous energy even higher at the bounce point, we refer to $\rho_{s\downarrow}^{1/4}$ as the quasi-highest energy scale of the bouncing universe.
@@
Figure 1: Evolution of the comoving Hubble radius $|aH|^{-1}$ in (left) inflationary and (right) bouncing cosmologies Cheung:2016vze ; Li:2024dce ; Li:2025ilc . In the bounce the horizon shrinks in Phases I and III and grows in Phases II and IV; primordial GWs exit in I, reenter in II, exit again in III, and finally reenter in IV, producing a double horizon–crossing history Li:2024dce . Inflation maintains an almost constant “inflationary” energy scale, tightly constrained by current GW limits Caprini:2018mtu . By contrast, bouncing models attain their maximum (“bounce”) energy scale at the central bounce Li:2024dce ; Li:2025ilc , which we constrain here using PTA data.
@@
In a recent work Li:2024dce we established a general framework for the generation and evolution of primordial gravitational waves (PGWs) in non-singular bouncing cosmologies. We partition the cosmic history into four phases—contracting collapse (I), bouncing contraction (II), bouncing expansion (III), and post-bounce radiation (IV)—as illustrated in Fig. 1. Employing the GW–propagation–matrix method of Ref. Li:2024dce , we derived an analytical expression for the PGW spectrum in bouncing cosmologies (c.f. Eq. (53) of Ref. Li:2024dce ).
Building on this, Ref. Li:2025ilc specialized to a concrete bounce by imposing three assumptions: (i) Phase IV is radiation dominated ($w_{4}=1/3$), as in the standard post-reheating Universe; (ii) the immediate pre- and post-bounce phases are governed by an exotic fluid with $w_{2}=w_{3}\to-\infty$ (e.g. Quintom matter Cai:2009zp ), ensuring a non-singular bounce; and (iii) the bounce is symmetric in conformal time, $\eta_{s\downarrow}\equiv\eta_{1\downarrow}=\eta_{3\downarrow}$. Under these conditions one obtains the closed-form SGWB spectrum in Eq. (1) (c.f. Eq. (24) of Ref. Li:2025ilc ), which explicitly depends on the bounce energy scale $\rho_{s\downarrow}^{1/4}$ (Eq. (2)).
In this work we employ the explicit SGWB spectrum for a non-singular bounce (Eq. (1)) to place the first direct PTA constraints on the bounce energy scale $\rho_{s\downarrow}^{1/4}$ in a generic four-phase bouncing-cosmology framework. Fitting to the combined NANOGrav 15-yr, EPTA DR2, PPTA DR3 and IPTA DR2 datasets NANOGrav:2023gor ; EPTA:2023fyk ; Reardon:2023gzh ; Antoniadis:2022pcn ; ZenodoNG ; NANOGrav:2023hde ; ZenodoEPTA ; EPTA:2023sfo ; PPTADR3 ; Zic:2023gta ; IPTADR2 ; Perera:2019sca via full Bayesian inference, we uncover two distinct posterior branches in $(\rho_{s\downarrow}^{1/4},w_{1})$: one at $w_{1}\approx 0.3$ and one at $w_{1}\gg 1$. Recasting these bounds into the dual $(a_{s\downarrow},w_{1})$ parametrization yields fully consistent results.
@@
    Dominant-energy-condition violation. The “right” branch ($w_{1}\gg 1$) violates the dominant energy condition (Table 1), offering empirical impetus to explore novel pre-bounce physics such as ghost condensates, higher-derivative or modified-gravity operators, and extra-dimensional effects.
@@
    consistent with the lower bound found in Ref. Li:2025ilc (for the forecast constraints on $\rho_{s\downarrow}^{1/4}$ from high frequency gravitational wave using superconducting LC circuit and resonant cavities, see Ref. Li:2025mxj ). Thus, current PTAs are already probing trans-Planckian regimes in bouncing models. Any proposed UV completion must respect these observational constraints, providing a rare window into physics beyond the Planck scale.
@@
2 Stochastic Gravitational Waves for Bouncing Cosmologies
In Ref. Li:2025ilc , considering a realistic bouncing cosmology that reheats to the standard radiation dominated era ($w_{4}=1/3$) after a sharp, symmetric bounce ($w_{2}=w_{3}\to-\infty$, $\eta_{s\downarrow}\equiv\eta_{1\downarrow}=\eta_{3\downarrow}$), we derive the PGW-induced stochastic gravitational-wave background (SGWB) spectrum (c.f. Eq.(24) in Ref. Li:2025ilc ), which encodes the bounce energy scale $\rho_{s\downarrow}^{1/4}$:
$\begin{split}\Omega_{\mathrm{GW}}(f)\,h^{2}&=\frac{h^{2}}{24}\left(\frac{f_{H_{0}}}{f_{m_{\mathrm{pl}}}}\right)^{2}\frac{C(w_{1})}{(2\pi)^{-n_{T}(w_{1})-1}}\\ &\quad\times\left(\frac{f}{f_{H_{0}}}\right)^{n_{T}(w_{1})}\left[\frac{\rho_{s\downarrow}^{1/4}}{(\rho_{c0}/\Omega_{\gamma 0})^{1/4}}\right]^{4-n_{T}(w_{1})}\mathcal{T}_{\mathrm{eq}}(f)\,.\end{split}$ (1)
Here, $\rho_{s\downarrow}^{1/4}$ is defined at the beginning of Phase IV (the reheating point) with the assumption that Phase IV is standard post-reheating radiation-dominated era (c.f. Eq.(20) in Ref. Li:2025ilc ),
$\rho_{s\downarrow}\equiv\rho(\eta_{s})=3H^{2}(\eta_{s\downarrow})m_{\mathrm{pl}}^{2}\leavevmode\nobreak\ ,$ (2)
$C(w_{1})$ is the amplitude factor Li:2025ilc ,
@@ 2 Stochastic Gravitational Waves for Bouncing Cosmologies
$n_{T}(w_{1})$ is the tilt of the primordial tensor spectrum, depending on $w_{1}$ Li:2025ilc ,
$n_{T}\equiv\frac{\mathrm{d}\ln\mathcal{P}_{h}(\eta_{k})}{\mathrm{d}\ln k}=\begin{cases}3-2x(w_{1}),&-\tfrac{1}{3}\leq w_{1}<1.\\ 3+2x(w_{1}),&w_{1}\geq 1,\end{cases},$ (4)
where $w_{i}$ denotes the equation-of-state parameter in phase $i$, $x(w_{1})\equiv\tfrac{3(1-w_{1})}{2(3w_{1}+1)}$ and $\Gamma(z)=\int_{0}^{\infty}t^{z-1}e^{-t}\,\mathrm{d}t$ is the Gamma function. The present-day critical density is $\rho_{c0}=3H_{0}^{2}m_{\mathrm{pl}}^{2}=8.098\,h^{2}\times 10^{-11}\,\mathrm{eV}^{4}$ with the reduced Hubble parameter $h=0.677$. Reference frequencies $f_{H_{0}}=2.2\times 10^{-18}\,\mathrm{Hz}$ and $f_{m_{\mathrm{pl}}}=3.7\times 10^{42}\,\mathrm{Hz}$ correspond to $H_{0}$ and the reduced Planck mass $m_{\mathrm{pl}}\equiv(\sqrt{8\pi G})^{-1}$, respectively, in natural units ($\hbar=c=1$). The radiation density fraction today is $\Omega_{\gamma 0}h^{2}=2.474\times 10^{-5}$. The transfer function
@@
captures the evolution from radiation to matter domination, with $f_{\mathrm{eq}}=2.01\times 10^{-17}\,\mathrm{Hz}$ marking matter–radiation equality. For the frequency band accessible to present gravitational-wave experiments—CMB/BICEP, pulsar timing arrays, and Advanced LIGO–Virgo—the equality-era transfer function satisfies $\mathcal{T}_{\mathrm{eq}}(f)\simeq 1$ because $f\gg f_{\mathrm{eq}}$. Accordingly, the stochastic gravitational-wave background observed today inherits the primordial tensor tilt $n_{T}(w_{1})$.
Figure 3 shows the tensor spectral tilt $n_{T}(w_{1})$ (red curve) obtained from Eq. (4). The stochastic background is red-tilted ($n_{T}<0$) for $w_{1}<0$, blue-tilted ($n_{T}>0$) for $w_{1}>0$, and strictly scale-invariant ($n_{T}=0$) at $w_{1}=0$ (a matter-dominated bounce). Consequently, low-frequency detectors (CMB/BICEP)—most sensitive to red tilts—set the strongest bounds in the $w_{1}<0$ regime, whereas high-frequency instruments (aLIGO+aVirgo) constrain more tightly when $w_{1}>0$.
@@
Current pulsar-timing-array (PTA) measurements prefer a blue-tilted spectrum: the NANOGrav 15-year data yield $n_{T}=1.8\pm 0.3$ (68% C.L.) Vagnozzi:2023lwo , shown as the yellow band in Fig. 3. Its intersection with the red curve selects two allowed intervals,
@@
3 Interpretation of Pulsar-Timing-Array Data in Bouncing Cosmology
3.1 Posterior distribution of bounce energy scale and physics interpretation
To determine the region in the joint parameter space $(\rho_{s\downarrow}^{1/4},w_{1})$ of bouncing cosmologies that is compatible with current PTA measurements, we carry out a Bayesian fit of Eq. (1) to the full PTA dataset ( NANOGrav 15-year, EPTA DR2, PPTA DR3 and IPTA DR2) NANOGrav:2023gor ; EPTA:2023fyk ; Reardon:2023gzh ; Antoniadis:2022pcn ; ZenodoNG ; NANOGrav:2023hde ; ZenodoEPTA ; EPTA:2023sfo ; PPTADR3 ; Zic:2023gta ; IPTADR2 ; Perera:2019sca , employing Ceffyl Lamb:2023jls . In Fig. 4 we present the resulting two-dimensional posterior distribution in the $(\rho_{s\downarrow}^{1/4},w_{1})$ plane for the bouncing scenario. Fig. 4 shows two distinct posterior branches, one near $w_{1}\approx 0.3$ and the other at $w_{1}\gg 1$, consistent with our analytic behavior of $n(w_{1})$ in Eq. (6). The small offset between the posterior peaks and the analytic prediction ($0.370.20<w_{1}<$ 0.37 and $w_{1}>13$) arises from the statistical uncertainties in the PTA data. More precisely, Fig. 4 shows that on the “left” branch the inferred bounce scale $\rho_{s\downarrow}^{1/4}$ grows with $w_{1}$ (see the high-resolution inset in Fig. 5), reflecting the monotonic rise of $n_{T}(w_{1})$ in that region, whereas on the “right” branch $\rho_{s\downarrow}^{1/4}$ asymptotes to a constant value as $n_{T}(w_{1})\to 2$ for $w_{1}\gg 1$ (the amplitude factor $C(w_{1})$ must also be accounted for in a more detailed description). Additionally, in Fig. 6, we show violin plots for two representative points—one from each of the $1\sigma$ posterior branches in the $(\rho_{s\downarrow}^{1/4},w_{1})$ plane (Fig. 4)—alongside examples in the $(a_{s\downarrow},w_{1})$ parametrization (Fig. 8) and for six conventional SGWB source models (Sec. A). These plots confirm that both bounce-model branches identified in this work can explain the nanohertz SGWB observed by current PTA datasets.